The mass of an oxygen atom is half that of a sulphur atom.can wee deside on this basis the density of sulphur vapour relative to oxygen is 2

Grade:

4 Answers

Abhinav Shresth
11 years ago

by using the formula that
vapour density = (M Wt.)/2

we can comment that vapour density of sulphur  is 2 relative to osygen.

but remember that formula used above is an approximation. 

Shivam Dimri
11 years ago

not the density,

the VAPOUR DENSITY IS TWO!!

thers a difference b/w he two...

Vikas TU
11 years ago

Vapour density = molar mass/2
thus V.D OF O atom = 16/2

                             = 8

V.D of S = 32 / 2     = 16

 V.D OF SULPHUR/V.D OF OXYGEN ATOM = 16/8

                                                         = 2 

FITJEE
11 years ago

Yes  we can judge it like that if it is molecular weight/ atomic weight--- since the formula of the vapour pressure is ------- (molec.. weigt)/2 

and  of O is 16 and S is 32.